OpenStudy (anonymous):

How do I factor 6x + 6y

OpenStudy (anonymous):

\[6x + 6y\] \[6( x + y)\]

OpenStudy (anonymous):

yeah that is the only way i can think of.

geerky42 (geerky42):

Yeah, this is only way to factor it.

OpenStudy (anonymous):

How do I factor 4y + 28 + 12z then? what formula do i use

geerky42 (geerky42):

dax + dby + dcz = d(ax+by+cz)

OpenStudy (anonymous):

\[4y + 28 + 12z\] \[4 (y + 7 + 3z)\]

OpenStudy (anonymous):

technically it would be more accurate if you wrote it like this \[4 (y+3 z+7)\]

OpenStudy (anonymous):

What happened to the 28 and the 12?

geerky42 (geerky42):

were divided by 4

geerky42 (geerky42):

28 = 4·7 12 = 4·3

OpenStudy (anonymous):

you see how the 28 and 12 have the 4 in common?

OpenStudy (anonymous):

I see so for example 2x + 8y + 16d would be 2(x + 4y + 8d)?

OpenStudy (anonymous):

\[2x + 8y + 16d\] \[2 (8 d+x+4 y)\]

OpenStudy (anonymous):

so yeah @JamesR4494 you are correct :)

OpenStudy (anonymous):

Aright what about 5x + 9d + 16c? how would that go

geerky42 (geerky42):

No common; this cannot be factored.

OpenStudy (anonymous):

\[5x + 9d + 16c\] Prime

OpenStudy (anonymous):

when it cant be factored, we say its PRIME.

OpenStudy (anonymous):

Oh okay.

OpenStudy (anonymous):

Thank you all for the help.

OpenStudy (anonymous):

yw :)

geerky42 (geerky42):